OpenStudy (anonymous):

tan^2 theta= (-3/2)sec theta (what values satisfy the equation?)

OpenStudy (cwrw238):

use the identity tan^2 theta = sec^2 theta - 1

OpenStudy (anonymous):

I dont have a calculator with me... how else can I solve this problem? I am so confused!

OpenStudy (jdoe0001):

hehe, so you have a computer but not a calculator :/

OpenStudy (cwrw238):

plug this in to your equation yo get an equation in sec theta only then solve for sec theta

OpenStudy (jdoe0001):

that's almost like saying I don't have a car just a ferrari

OpenStudy (cwrw238):

you'll have a calculator on your computer

OpenStudy (jdoe0001):

$$ tan^2(\theta)= -\cfrac{3}{2}sec(\theta)\\ \color{red}{1+tan^2(\theta) = sec(\theta)^2 \implies tan^2(\theta) = sec(\theta)^2-1}\\ sec(\theta)^2-1= -\cfrac{3}{2}sec(\theta)\\ 2sec(\theta)^2-2= -3sec(\theta)\\ 2sec(\theta)^2+3sec(\theta)-2=0\\ $$